A sofa purchased 5 years ago for $2,000 has been destroyed by a tornado.  It is 60% depreciated. Today the same sofa costs $2,250. If a homeowner files a claim for the lost sofa, how much would the insurer pay, assuming there is no deductible?
Answer #1

Answer:

Cost of Same Sofa today S=2250

Depreciation D =60%

Actual cash basis

Insurer will pay = S*(1-D)=2250*(1-60%)=$900